OpenStudy (crystal_bliss):

I don't get this question or what it asking can anyone plz help.... Line Q is represented by the following equation: 2x + y = 11 Which equation completes the system that is satisfied by the solution (3, 5)? A. x + 2y = 15 B. x + y = 15 C. x + y = 8 D. x – y = 2

OpenStudy (crystal_bliss):

not really

OpenStudy (crystal_bliss):

sure

OpenStudy (crystal_bliss):

11/2 is 5.5

OpenStudy (crystal_bliss):

not realy how would it be A?

OpenStudy (crystal_bliss):

ok i will se if it is A and i will tell you if you get it right.

OpenStudy (crystal_bliss):

@TheCuteOne25 it was wrong and i asked for help in the first place not wrong answers so i won't medal you

OpenStudy (phi):

When you are given two equations (of lines) the "solution" is the point where the lines meet (cross over each other) in other words, in this case, they say the point (3, 5) is the solution, so that point is on both lines. The first equation is 2x + y = 11 and if we put in 3 for x 2*3 + y= 11 and 5 for y 2*3+5 = 11 we see the left side simplifies to 6+5 or 11 and the equation is true. Now we need another equation where (3,5) "works" we should test each of the choices. Do you know how to do that ?

OpenStudy (crystal_bliss):

um not really

OpenStudy (phi):

A. x + 2y = 15 erase the x and put in 3 erase the y and put in 5 (and remember 2y really means 2*y)

OpenStudy (phi):

can you do that ?

OpenStudy (crystal_bliss):

so would it be 3+2*5=15 ?

OpenStudy (phi):

yes. now simplify the left side. what do you get ?

OpenStudy (crystal_bliss):

it would be 13=15

OpenStudy (phi):

and that is not true. 13 is not the same as 15. that means the point (3,5) does not "work" i.e. is not on the line but we want an equation where (3,5) is on the line. Try the next equation: B. x + y = 15

OpenStudy (phi):

replace the x with 3 and replace the y with 5 then simplify the left side what do you get?

OpenStudy (phi):

you are almost to the answer. we are checking if (3,5) works for equation B

OpenStudy (crystal_bliss):

i don't think it would be the next line either i think it would be C the third line because it makes a true equation which is 5+3=8 8=8

OpenStudy (crystal_bliss):

would that be right?

OpenStudy (phi):

yes, C C. x + y = 8 works. (3,5) is on that line and (3,5) is on the first line 2x + y = 11 if a point is on two lines at the same time, that is where the lines cross. and that is the solution of the system (as they say) anyway, C is the answer

OpenStudy (phi):

and to be complete, if we check D. x – y = 2 we find 3-5=2 -2 = 2 does not work (which is good, we only want one choice)

OpenStudy (crystal_bliss):

ok thank you soooo much for explaining it to me!! you are very helpfull

OpenStudy (phi):

yw